OpenStudy (anonymous):

-6p+7=3(2p-3)-4(-10+4p)

OpenStudy (anonymous):

-6p+7=3(2p-3)-4(-10+4p) First distribute -6p+7=6p-9+40-16p Combine like terms -6p+7=-10p+31 Add 10p and Subtract 7 on both sides 4p=24 divide by 4 on both sides p=6

OpenStudy (anonymous):

ok thanks